Huron Company produces a commercial cleaning compound known as Zoom. The direct materials and direct labor standards for oneWhat's the material quantity variance?
Answer #1

Solution:

Material price variance $1,160 F
Material quantity variance $10,440 U

Explanation:

Material price variance = AQ (AP - SP) = 11600(2.90-3) = 1,160 F

Material quantity variance = SP(AQ - SQ) = 3 (10440-6960*) = 10440 U

*1160 X 6 = 6960